OpenStudy (anonymous):

15x^7/9y^5 * 45y^2/5x^3 AND 15x^7/9y^5 ÷ 45y^2/5x^3

OpenStudy (anonymous):

\[15x^7/9y^5 * 45y^2/5x^3=15x ^{4}/y ^{2}\]

OpenStudy (anonymous):

\[15x^7/9y^5 ÷ 45y^2/5x^3=(15x^7*5x^3)/(9y^5* 45y^2)=5x ^{10}/27y^7\]

OpenStudy (anonymous):

Just use your rules for exponents.

OpenStudy (anonymous):

thanks